PLEASE HELP ME SOLVE THIS EQUATION: Te sum of two numbers is 8.5 and their product is 18. Find the numbers?

Thank you!!

Answer by devjyotsingh24(4)

Let one of the numbers be x.Then clearly the other no.=(8.5-x).
Using the other hint in the problem we get:
x*(8.5-x)=18
Thus x^x-8.5x+18=0
Now this is a pretty straightforward quadratic equation which should be solved using the quadratic formula.